OpenStudy (anonymous):

A nylon rope used by mountaineers elongates 1.50m under the weight of a 66.0kg climber.If the rope is 48.5m in length and 6.90mm in diameter, what is Young's modulus for this nylon?

OpenStudy (anonymous):

First thing I did was to convert the diameter into meters =0.0069,r=0.00345 cross area would be (3.14)(r)^2= 3.74*10^-5, E=Stress/Strain. The Stress=F/A, and Strain \[DeltaL\]/lo

OpenStudy (anonymous):

My numbers for stress are: 646.8/3.71*10^-5=1.73*10^.3. For Strain-1.5/48.5=3.09*10^-2, So my Y=5.59*10^-6, but apparently this answer is not correct :( can someone tell me what I did wrong?

OpenStudy (anonymous):

I got it :) little mistake again :)

OpenStudy (anonymous):

answer is 5.56*10^8

OpenStudy (anonymous):

lol k.